OpenStudy (anonymous):

Use identities (no calculators) to fi nd the exact value for cos(pi/2)cos(pi/6) - sin(pi/2)sin(pi/6)

OpenStudy (anonymous):

ho ho ho this one is \[cos(a+b)\] \[a=\frac{\pi}{2}, b=\frac{\pi}{6}, a+b=\frac{2\pi}{3}\] \[cos(\frac{\pi}{3})=\frac{1}{2}\]